what is the surface area of a cube that has sides of a length of 4 cm
Found 3 solutions by rfer, Alan3354, mananth:
Answer by rfer(16322)

4*4*6=96 sq cm

Answer by Alan3354(69443)

A cube has 6 sides, all 6 are squares.
Find the area of a square with sides of 4 cm, multiply that by 6.

Answer by mananth(16946)

If you observe there are six faces in a cube.
The area of each face = 4*4 = 16sq.cm.
There are 6 faces so
total surface area = 6* 16 = 96 sq. cm